# A chaotic lattice field theory in two dimensions

### v2

@@ Appendix B Computation of spatiotemporal cat periodic states
Defining equations Eq. (49) are piecewise linear, and, given a primitive cell $\mathbb{A}$ and a mosaic $\mathsf{M}$ Eq. (37) over it, always has a unique solution $\mathsf{\Phi}_{\mathsf{M}}$}. We solve it by reciprocal lattice diagonalization (Sec. VIII.2), by direct determinant evaluation (Appendix B.3), or by matrix inversion:
@@
where $g_{zz^{\prime}}$, the inverse of the orbit Jacobian operator, is the Klein-Gordon free-field Eq. (48) Green’s function. In literature, $g_{zz^{\prime}}$ is known as the Green’s function for the $d$-dimensional discretized screened Poisson equation.
@@ B.1 Spatiotemporal cat mosaics
We have no algorithm that would generate admissible spatiotemporal cat mosaics (see question 6 of Sec. XII.2 Open questions). Instead, we solve the linear equation Eq. (161) for each covering mosaic, and then -for mosaics containing exterior alphabets $\mathcal{A}_{1}$ letters- discard those for which $\mathsf{\Phi}_{\mathsf{M}}$} lies outside the unit hypercube Eq. (162).
@@ B.2 Spatiotemporal cat primitive cells’ orbit Jacobians
(Continuation of calculations of Sec. VIII.) Developing some feel for the orbit Jacobian formulas for two-dimensional spatiotemporal cat examples is now in order. The simplest examples of periodic states, illustrated by spatiotemporal mosaic tilings of Fig. 6, are (i) spacetime steady states over the unit cell $1\!\times\!1]_{0}$}, (ii) spatial steady states over $1\!\times\!{T}]_{0}$}, (iii) temporal steady states over ${L}\!\times\!1]_{0}$}, and (iv) time-relative steady states over ${L}\!\times\!1]_{{S}}$}, ${S}\neq 0$, stationary patterns in a time-reference frame [216] moving with a constant velocity S/T{S}/{T}.
@@
For a list of such two-dimensional spatiotemporal cat orbit Jacobians, see Table 1, and the list of the spatiotemporal cat orbit Jacobians evaluated for $\mu^{2}=1$, see Table 2.
@@ B.3 Spatiotemporal cat: Fundamental fact
We now show that for a spatiotemporal cat orbit Jacobian counts the number of periodic states in any spatiotemporal dimension $d$.
@@
means geometrically. The ${N}_{\mathbb{A}}\!\times\!{N}_{\mathbb{A}}]$ orbit Jacobian matrix ${\mathcal{J}}_{\mathbb{A}}$ stretches the state space unit hypercube $\mathsf{\Phi}\in[0,1)^{{N}_{\mathbb{A}}}$ into an ${N}_{\mathbb{A}}$-dimensional fundamental parallelepiped (or parallelogram), and maps the periodic state ${\mathsf{\Phi}}_{\mathsf{M}}$ into a point on integer lattice $\mathbb{Z}^{{N}_{\mathbb{A}}}$ within it, in the ${N}_{\mathbb{A}}$-dimensional configuration state space Eq. (12). This point is then translated by integer winding numbers $\mathsf{M}$ into the origin. What Baake et al. [217] call the ‘fundamental fact’ follows: margin: [Uncaptioned image] margin: [Uncaptioned image]
@@
For the history of ‘fundamental fact’ see Appendix A. Historical context of the companion paper I. [16] Reader might also want to check the figures of a few fundamental parallelepipeds there, but we know of no good way of presenting them visually for primitive cells of interest here, with ${N}_{\mathbb{A}}>3$.
It is a peculiarity of the spatiotemporal cat that it involves two distinct integer lattices. (i) The spacetime coordinates Eq. (6) are discretized by integer lattice $\mathbb{Z}^{d}$. The primitive cell $\mathbb{A}$ Eq. (10) is an example of a fundamental parallelepiped, and we use the fundamental fact when we express the volume Eq. (11) of the primitive cell, i.e. the determinant of the matrix $\mathbb{A}$, as the number of lattice sites within the primitive cell. (ii) For a spatiotemporal cat the lattice site field $\phi_{z}$ Eq. (48) is compactified to the unit circle $[0,1)$, imparting integer lattice structure to the configuration state space Eq. (12): the orbit Jacobian matrix ${\mathcal{J}}_{\mathbb{A}}$ maps a periodic state $\mathsf{\Phi}_{\mathsf{M}}\in[0,1)^{{N}_{\mathbb{A}}}$ to a $\mathbb{Z}^{{N}_{\mathbb{A}}}$ integer lattice site $\mathsf{M}$. Nothing like that, and no ‘fundamental fact’ applies to general nonlinear field theories of Sec. III.
@@
As a concrete example consider periodic states of two-dimensional spatiotemporal cat with periodicity $3\!\times\!2]_{0}$}, i.e., space period L=3{L}=3, time period T=2{T}=2 and tilt S=0{S}=0. Periodic states within the primitive cell and their corresponding mosaics can be written as two-dimensional $3\!\times\!2]$ arrays:
@@
Reshape the periodic states and mosaics into vectors:
@@
The reshaped orbit Jacobian matrix acting on these periodic states is a block matrix:
@@
where the stretching factor $2s=4+\mu^{2}$. The fundamental parallelepiped generated by the action of orbit Jacobian matrix ${\mathcal{J}}_{{[3\!\times\!2]_{0}}}$ on the state space unit hypercube Eq. (48) is spanned by 6 primitive vectors, the columns of the orbit Jacobian matrix Eq. (181). The ‘fundamental fact’ now expresses the orbit Jacobian, i.e., the number of periodic states within the fundamental parallelepiped, as a polynomial of order ${N}_{\mathbb{A}}$ in the Klein-Gordon mass $\mu^{2}$ Eq. (60),
@@
without recourse to any explicit diagonalization, such as the reciprocal lattice diagonalization Eq. (93). For $\mu^{2}=1$ this agrees with the reciprocal lattice evaluation Eq. (170). For a list of the numbers of spatiotemporal cat periodic states for primitive cells ${L}\!\times\!{T}]_{S}$} up to $3\!\times\!3]_{2}$}, see Table 1.
For $\mu^{2}=1$ spatiotemporal cat the pruning turns out to be very severe. Only 52 of the prime $2\!\times\!2]_{0}$} mosaics are admissible. As for the repeats of smaller mosaics, there are 2 admissible $1\!\times\!2]_{0}$} mosaics repeating in time and 2 $2\!\times\!1]_{0}$} mosaics repeating in space. There are 4 admissible $1/2$-shift periodic boundary $1\!\times\!2]_{0}$} mosaics. And there is 1 admissible mosaic which is a repeat of letter 0. The total number of $2\!\times\!2]_{0}$} of periodic states is obtained by all cyclic permutations of admissible prime mosaics,
@@ B.4 Prime lattice field configurations
The enumeration of spatiotemporal cat doubly-periodic states proceeds in 3 steps:
@@
    We have defined the prime orbit in Sec. VI.
@@
Given the number of periodic states, the number of $\mathbb{A}={[{L}\!\times\!{T}]_{S}}$-periodic prime orbits is computed recursively:
@@ B.5 Example: ‘Escape rate’ of temporal cat
The topological zeta function of temporal cat is: [218, 16]
@@
where ${N}_{{n}}$ is the number of periodic states with period n{n}. Due to the uniform stretching factor $\mu^{2}+2$, the deterministic zeta function of temporal cat has the same form, up to a rescaling:
@@
The leading root is 1 so the ‘escape rate’ is 0. The Fredholm determinant [152] of temporal cat is:
@@
where we have used the ‘fundamental fact’ Eq. (172). The ‘escape rate’ is again 0, as it should be - cat map is by construction probability conserving.
B.6 Example: Expectation value of stability exponent of temporal cat.
To compute the expectation value of the stability exponent, take the logarithm of periodic state’s primitive cell stability as the Birkhoff sum $A$, Eq. (19), stability exponent observable, and compute the corresponding deterministic zeta function:
@@
where $|\mathrm{Det}\,{\mathcal{J}}_{n}|$ is the primitive cell stability of period-$n$ periodic states, and $\Lambda$ is the stability multiplier which is related to the stability exponent by Eq. (187). Note that the number of $n$-periodic state is given by the primitive cell stability: [218, 16]
@@
Using Eq. (137) the expectation value of the stability exponent is:
@@
which agrees with the fact that every periodic state has a same stability exponent $\lambda$.
@@ Appendix C Spectra of orbit Jacobian operators for nonlinear field theories
The simplicity of the spatiotemporal cat orbit Jacobian operator band spectrum Eq. (106), plotted in Fig. 9 (a) and Fig. 11 (a), is a bit misleading. As explained in Sec. IV.2, the uniform stretching factor describes only the stability of a steady state solution, for any field theory. To get a feeling for the general case, in section 10 of paper I [16] we compute the stability of a period-2 periodic state for two nonlinear field theories. Here we outline such calculations, to illustrate the essential difference between the very special spatiotemporal cat case, and the general, nonlinear case. For a detailed exposition, see companion paper III, [17] where we evaluate stabilities of large sets of nonlinear field theories’ periodic states.
@@ C.1 One-dimensional $\phi^{3}$ field theory period-2 periodic state
plotted in Fig. 9 (b), in the $k\in(-\pi/2,\pi/2]$ Brillouin zone for $\mu^{2}=3.5$. For a finite primitive cell of even period, tiled by $r$th repeat of the period-2 periodic state $\mathsf{\Phi}_{p}$, the eigenvalues of its orbit Jacobian matrix are $\Lambda_{01,\pm}(k)$ evaluated at $k$ restricted to a discrete set of wave vectors $k$, multiples of $\pi/r$: an example is worked out in Sec. VIII.3, with third and fourth repeats plotted in Fig. 9 (b).
@@ C.2 Two-dimensional $\phi^{4}$ field theory $2\!\times\!1]_{0}$ periodic state
to which companion paper III [17] assigns alphabet Eq. (36) $\mathcal{A}=\{-{1},0,1\}\,$, has at most 3 steady states. The two spacetime dimensions $\phi^{4}$ has has at most 3 period-2 prime orbits of periodicity $2\!\times\!1]_{0}$, with mosaics $\begin{array}[]{|c|c|}\hline\cr-1&0\cr\hline\cr\end{array}\,$, $\begin{array}[]{|c|c|}\hline\cr-1&1\cr\hline\cr\end{array}$ and $\begin{array}[]{|c|c|}\hline\cr 0&1\cr\hline\cr\end{array}\,.$ For example, for Klein-Gordon mass-squared $\mu^{2}=5$, one of the period-2 prime orbits is
@@
The orbit Jacobian operator has two Bloch bands:
@@
plotted in Fig. 11 (b). While the period-2 periodic state $2\!\times\!1]_{0}$ is the same in one and two spatiotemporal dimensions, it stability, Eq. (199), is evaluated in two dimensions, with transverse $k_{2}$, temporal direction eigenstates included. For any finite primitive cell tiled by repeats of the prime orbit $\mathsf{\Phi}_{p}$, eigenstates of the orbit Jacobian matrix have a discrete set of wave vectors $k$. As an example, eigenvalues of a $6\!\times\!4]_{0}$ periodic state tiled by 12 repeats of $\mathsf{\Phi}_{p}$ have wave vectors $k$ marked by black dots in Fig. 11 (b).

### v3

@@ Appendix B Computation of spatiotemporal cat periodic states
Defining equations Eq. (49) are piecewise linear, and, given a primitive cell $\mathbb{A}$ and a mosaic $\mathsf{M}$ Eq. (37) over it, always has a unique solution $\mathsf{\Phi}_{\mathsf{M}}$}. We solve it by reciprocal lattice diagonalization (Sec. IX.2), by direct determinant evaluation (Appendix B.3), or by matrix inversion,
@@
where $g_{zz^{\prime}}$, the inverse of the orbit Jacobian operator, is the Klein-Gordon free-field Eq. (48) Green’s function. In the literature, $g_{zz^{\prime}}$ is known as the Green’s function for the $d$-dimensional discretized screened Poisson equation.
@@ B.1 Spatiotemporal cat mosaics
We have no algorithm that would generate admissible spatiotemporal cat mosaics (see question 6 of Sec. XIII.2). Instead, we solve the linear equation Eq. (161) for each covering mosaic, and then -for mosaics containing exterior alphabets $\mathcal{A}_{1}$ letters- discard those for which $\mathsf{\Phi}_{\mathsf{M}}$} lies outside the unit hypercube Eq. (162).
@@ B.2 Spatiotemporal cat primitive cells’ orbit Jacobians
This section provides a continuation of calculations of Sec. IX. Developing some feel for the orbit Jacobian formulas for two-dimensional spatiotemporal cat examples is now in order. The simplest examples of periodic states, illustrated by spatiotemporal mosaic tilings of Fig. 4, are (i) spacetime steady states over the unit cell $1\!\times\!1]_{0}$}, (ii) spatial steady states over $1\!\times\!{T}]_{0}$}, (iii) temporal steady states over ${L}\!\times\!1]_{0}$}, and (iv) time-relative steady states over ${L}\!\times\!1]_{{S}}$}, ${S}\neq 0$, stationary patterns in a time-reference frame [229] moving with a constant velocity S/T{S}/{T}.
@@
For a list of such two-dimensional spatiotemporal cat orbit Jacobians, see Table 1, and for a list of the spatiotemporal cat orbit Jacobians evaluated for $\mu^{2}=1$, see Table 2.
@@ B.3 Spatiotemporal cat: Fundamental fact
We now show that for a spatiotemporal cat, orbit Jacobian counts the number of periodic states in any spatiotemporal dimension $d$.
@@
means geometrically. The ${N}_{\mathbb{A}}\!\times\!{N}_{\mathbb{A}}]$ orbit Jacobian matrix ${\mathcal{J}}_{\mathbb{A}}$ stretches the state space unit hypercube $\mathsf{\Phi}\in[0,1)^{{N}_{\mathbb{A}}}$ into an ${N}_{\mathbb{A}}$-dimensional fundamental parallelepiped (or a parallelogram), and maps the periodic state ${\mathsf{\Phi}}_{\mathsf{M}}$ into a point on integer lattice $\mathbb{Z}^{{N}_{\mathbb{A}}}$ within it, in the ${N}_{\mathbb{A}}$-dimensional configuration state space Eq. (12). This point is then translated by integer winding numbers $\mathsf{M}$ into the origin. What Baake et al. [230] call the ‘fundamental fact’ follows: [231] margin: [Uncaptioned image] margin: [Uncaptioned image]
@@
For the history of ‘fundamental fact’ see Appendix A. Historical context of the companion paper I. [16] The reader might also want to check the figures of a few fundamental parallelepipeds there, but we know of no good way of presenting them visually for primitive cells of interest here, with ${N}_{\mathbb{A}}>3$.
It is a peculiarity of the spatiotemporal cat that it involves two distinct integer lattices. (i) The spacetime coordinates Eq. (6) are discretized by integer lattice $\mathbb{Z}^{d}$. The primitive cell $\mathbb{A}$ Eq. (10) is an example of a fundamental parallelepiped, and we use the fundamental fact when we express the volume Eq. (11) of the primitive cell, i.e. the determinant of the matrix $\mathbb{A}$, as the number of lattice sites within the primitive cell. (ii) For a spatiotemporal cat, the lattice site field $\phi_{z}$ Eq. (48) is compactified to the unit circle $[0,1)$, imparting an integer lattice structure to the configuration state space Eq. (12): the orbit Jacobian matrix ${\mathcal{J}}_{\mathbb{A}}$ maps a periodic state $\mathsf{\Phi}_{\mathsf{M}}\in[0,1)^{{N}_{\mathbb{A}}}$ to a $\mathbb{Z}^{{N}_{\mathbb{A}}}$ integer lattice site $\mathsf{M}$. Nothing like that, and no ‘fundamental fact’ applies to general nonlinear field theories of Sec. IV.
@@
As a concrete example, consider periodic states of two-dimensional spatiotemporal cat with periodicity $3\!\times\!2]_{0}$}, i.e., space period L=3{L}=3, time period T=2{T}=2, and tilt S=0{S}=0. Periodic states within the primitive cell and their corresponding mosaics can be written as two-dimensional $3\!\times\!2]$ arrays,
@@
We reshape the periodic states and mosaics into vectors,
@@
The reshaped orbit Jacobian matrix acting on these periodic states is a block matrix,
@@
where the stretching factor $2s=4+\mu^{2}$. The fundamental parallelepiped generated by the action of orbit Jacobian matrix ${\mathcal{J}}_{{[3\!\times\!2]_{0}}}$ on the state space unit hypercube Eq. (48) is spanned by six primitive vectors, the columns of the orbit Jacobian matrix Eq. (181). The ‘fundamental fact’ now expresses the orbit Jacobian, i.e., the number of periodic states within the fundamental parallelepiped, as a polynomial of order ${N}_{\mathbb{A}}$ in the Klein-Gordon mass $\mu^{2}$ Eq. (60),
@@
without recourse to any explicit diagonalization, such as the reciprocal lattice diagonalization Eq. (93). For $\mu^{2}=1$, this agrees with the reciprocal lattice evaluation Eq. (170). For a list of the numbers of spatiotemporal cat periodic states for primitive cells ${L}\!\times\!{T}]_{S}$} up to $3\!\times\!3]_{2}$}, see Table 1.
For the $\mu^{2}=1$ spatiotemporal cat the pruning turns out to be very severe. Only 52 of the prime $2\!\times\!2]_{0}$} mosaics are admissible. As for the repeats of smaller mosaics, there are 2 admissible $1\!\times\!2]_{0}$} mosaics repeating in time and 2 $2\!\times\!1]_{0}$} mosaics repeating in space. There are 4 admissible $1/2$-shift periodic boundary $1\!\times\!2]_{0}$} mosaics. And there is 1 admissible mosaic which is a repeat of letter 0. The total number of $2\!\times\!2]_{0}$} of periodic states is obtained by all cyclic permutations of admissible prime mosaics,
@@ B.4 Prime lattice field configurations
The enumeration of spatiotemporal cat doubly periodic states proceeds in three steps:
@@
    We have defined the prime orbit in Sec. VII.
@@
Given the number of periodic states, the number of $\mathbb{A}={[{L}\!\times\!{T}]_{S}}$-periodic prime orbits is computed recursively,
@@ B.5 Example: ‘Escape rate’ of temporal cat
The topological zeta function of temporal cat is, [232, 16]
@@
where ${N}_{{n}}$ is the number of periodic states with period n{n}. Due to the uniform stretching factor $\mu^{2}+2$, the deterministic zeta function of temporal cat has the same form, up to rescaling,
@@
The leading root is 1, so the ‘escape rate’ is 0. The Fredholm determinant [159] of temporal cat is
@@
where we have used the ‘fundamental fact’ Eq. (172). The ‘escape rate’ is again 0, as it should be–the cat map is by construction probability conserving.
B.6 Example: Expectation value of the stability exponent of a temporal cat.
To compute the expectation value of the stability exponent, take the logarithm of periodic state’s primitive cell stability as the Birkhoff sum $A$, Eq. (19), stability exponent observable, and compute the corresponding deterministic zeta function,
@@
where $|\mathrm{Det}\,{\mathcal{J}}_{n}|$ is the primitive cell stability of period-$n$ periodic states, and $\Lambda$ is the stability multiplier, which is related to the stability exponent by Eq. (187). Note that the number of $n$-periodic states is given by the primitive cell stability, [232, 16]
@@
Using Eq. (137), the expectation value of the stability exponent is
@@
which agrees with the fact that every periodic state has the same stability exponent $\lambda$.
@@ Appendix C Spectra of orbit Jacobian operators for nonlinear field theories
The simplicity of the spatiotemporal cat orbit Jacobian operator band spectrum Eq. (106), plotted in Fig. 11 (a) and Fig. 10 (a), is a bit misleading. As explained in Sec. V.2, the uniform stretching factor describes only the stability of a steady state solution, for any field theory. To get a feeling for the general case, in Sect. 10 of paper I [16] we compute the stability of a period-2 periodic state for two nonlinear field theories. Here we outline such calculations, to illustrate the essential difference between the very special spatiotemporal cat case, and the general, nonlinear case. For a detailed exposition, see companion paper III, [17] where we evaluate stabilities of large sets of nonlinear field theories’ periodic states.
@@ C.1 One-dimensional $\phi^{3}$ field theory period-2 periodic state
plotted in Fig. 11 (b), in the $k\in(-\pi/2,\pi/2]$ Brillouin zone for $\mu^{2}=3.5$. For a finite primitive cell of even period, tiled by $r$th repeat of the period-2 periodic state $\mathsf{\Phi}_{p}$, the eigenvalues of its orbit Jacobian matrix are $\Lambda_{01,\pm}(k)$ evaluated at $k$ restricted to a discrete set of wave vectors $k$, multiples of $\pi/r$: an example is worked out in Sec. IX.3, with third and fourth repeats plotted in Fig. 11 (b).
@@ C.2 Two-dimensional $\phi^{4}$ field theory $2\!\times\!1]_{0}$ periodic state
to which companion paper III [17] assigns alphabet Eq. (36) $\mathcal{A}=\{-{1},0,1\}\,$, has at most three steady states. The two spacetime dimensions $\phi^{4}$ has at most three period-2 prime orbits of periodicity $2\!\times\!1]_{0}$, with mosaics $\begin{array}[]{|c|c|}\hline\cr-1&0\cr\hline\cr\end{array}\,$, $\begin{array}[]{|c|c|}\hline\cr-1&1\cr\hline\cr\end{array}\,$, and $\begin{array}[]{|c|c|}\hline\cr 0&1\cr\hline\cr\end{array}\,.$ For example, for Klein-Gordon mass-squared $\mu^{2}=5$, one of the period-2 prime orbits is
@@
The orbit Jacobian operator has two Bloch bands,
@@
plotted in Fig. 10 (b). While the period-2 periodic state $2\!\times\!1]_{0}$ is the same in one and two spatiotemporal dimensions, its stability, Eq. (199), is evaluated in two dimensions, with transverse $k_{2}$, temporal direction eigenstates included. For any finite primitive cell tiled by repeats of the prime orbit $\mathsf{\Phi}_{p}$, eigenstates of the orbit Jacobian matrix have a discrete set of wave vectors $k$. As an example, eigenvalues of a $6\!\times\!4]_{0}$ periodic state tiled by 12 repeats of $\mathsf{\Phi}_{p}$ have wave vectors $k$ marked by black dots in Fig. 10 (b).